OpenStudy (anonymous):

Select all of the steps that could be used to solve the equation 1 + 3x = -x + 4. add x, subtract 1, divide by 4 add x, subtract 4, divide by 4 subtract 3x, subtract 4, divide by 4 subtract 3x, subtract 4, divide by -4 subtract 1, add x, divide by 4

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

1 + 3x = -x + 4 You can first add 'x' to both sides: 1 + 4x = 4 Now what's our next step?

OpenStudy (anonymous):

We want '4x' by itself